OpenStudy (anonymous):

simple harmonic motion , d=8cos((3pi)/(4)t) answer the following questions: Maximum: 8 Frequency: ????

OpenStudy (anonymous):

Let's recall. \[x(t) = A \sin(\omega t)\]where \(\omega\) is the angular velocity. Frequency can be expressed in terms of \(\omega\) as such\[f = {\omega \over 2 \pi}\]

OpenStudy (anonymous):

4/2pi (Is that correct?)

OpenStudy (anonymous):

I would imagine that \(\omega = 3 \pi /4\)

OpenStudy (anonymous):

Ok I got it, I was looking at the wrong thing THANK YOU :)

OpenStudy (anonymous):

What about this question: 0=8cos(3pi/4)(t) So far I got: 0=8(-√2/2)(t)

OpenStudy (anonymous):

What are you solving for with this equation?

OpenStudy (anonymous):

It says: Determine the least positive value of t which d=0

OpenStudy (anonymous):

You can just solve\[\cos({3 \pi \over 4} t ) = 0\]

OpenStudy (anonymous):

We know that \(\cos(\pi/2) = 0\) Therefore, \[{\pi \over 2} = {3 \pi \over 4} t\]